PuzzledGurl:

Yesterday, the snow was 2 feet deep in front of Archie’s house. Today, the snow depth dropped to 1.6 feet because the day is so warm. What is the percent change in the depth of the snow? I think it should be over 50% right?

Ultrilliam:

First, we need to subtract the second figure from the first figure. So \(1.6-2\) Then, we're going to want to divide that number by 2, and finally multiply that by 100. Our final problem should look a little like this: \(1.6-2=x\) \(x/2=y\) \(y*100=z\)

PuzzledGurl:

so 1.6-2 will be negative? It will = -0.4 -0.4/2=-0.2 and -0.2*100=-20?

Ultrilliam:

Correct, your percent change is 20%, and it was a 20% decrease

PuzzledGurl:

Okay

PuzzledGurl:

So if I had another question: Lydia bought a shirt at 20% off its retail price of $40. She paid 5% tax on the price after the discount. How much did Lydia pay for the shirt?, would I do the same thing?

Ultrilliam:

Could you actually ask that in a new question? (Close this one, and ask another one)

PuzzledGurl:

All I am asking is if I would do the same thng or not but ok

Ultrilliam:

Ah, no, it's a little different for this one